OpenStudy (anonymous):

A small object is thrown out of a hot air balloon from an unknown height. It falls straight down and hits the ground 5.6 seconds later with a speed of 125 m/s. How high is the balloon?

OpenStudy (shane_b):

Information given:\[V_f=125m/s\]\[a=9.8m/s^2\]\[t=5.6s\]Since you only have constant acceleration (due to gravity) you can use the kinematic equations (see image). First, find the initial velocity at which the object was thrown downward: \[\large V_f=V_i+at\]\[\large V_i=V_f-at=125m/s-(9.8m/s^2)(5.6s)=70.12m/s\] Now calculate the distance:\[\large V_f^2=V_i^2+2ad\]\[\large d=\frac{V_f^2-V_I^2}{2a}=546.336m\]

OpenStudy (anonymous):

I looked at this as a conservation of energy problem. That is:\[mgh=\frac{1}{2}mv^2\]The masses cancel out, leaving you with \[9.8\frac{m}{s^2}*h=\frac{1}{2}*(125\frac{m}{s})^2\]But I got:\[h=797m\]

OpenStudy (shane_b):

The problem is that the ball not only had PE at the top...it also had KE in the beginning since it was thrown downward.

OpenStudy (anonymous):

Got bumped. Crummy time for an update. :| Anyhow, yes, Shane_B, that makes sense. Doing this after your first step gives me the same answer:\[mgh + \frac{1}{2}m(70.12\frac{m}{s})^2=\frac{1}{2}m(125\frac{m}{s})^2\]

OpenStudy (shane_b):

Yea...I thought I broke OpenStudy again :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thank you guys so much :D